OpenStudy (anonymous):

f(x)=3x+2 to evaluate f(x)-f(5).... so f(x)-f(5)= ?

OpenStudy (anonymous):

\[f(5)=3\times 5-2=13\]

OpenStudy (anonymous):

so i guess \[f(x)-f(5)=3x+2-13=3x-11\]

OpenStudy (anonymous):

where is the -2 from? is it supposed to be +2?